AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME, Page 26 of 33  ELOG logo
  ID Date Author Subject
  157   Thu Nov 18 12:29:50 2021 CB RSSConditioning D2000 NEGs

D2000 thermocouples report temperatures around 100C (D2000 #1,2) and 120C (3,4). 1,2 are in the intermediate chamber, 3,4 are behind the moving arms. 3 is directly being the detectors.

Wanted to flash ion elements in D2000, but SAES recommend >100C to flash.

Followed SAES procedure / pre-sets to condition D2000 NEG elements while baking, but reduced temperature from 250 to 150 C (1,2) 140 C (4) and 130 C (3).

 

No appreciable changes in temperature. Change in pressure quite apparent. Left elements at temperatures above. May consider increasing everything to 150 tomorrow assuming detector temperatures do not change.

May also consider conditioning NEG elements in intermediate and DSSD wall chambers with 15W power (150C?) during baking as per SAES procedure. We never did this before, but we never baked as such low temperatures either. Will have to activate NEG elements after conditioning either way.

Attachment 1: vacplot.png
vacplot.png
Attachment 2: CARMEtemps.png
CARMEtemps.png
  156   Wed Nov 17 07:12:22 2021 RSS CB + ML (remotely)CARME baking with tape continues 17 Nov

CARME baking with heating tape continues

Left overnight at -

Ch1-3: 140 degrees

Ch5-6: 150 degrees

Ch8:    100 degrees

 

07:50 - Increased channels 1-3 to 150, channels 5 and 6 to 160, and channel 8 remain at 100. For the gas jet target, channel 2 increased to 150.

12:39 - Increased channels 5 and 6 to 170 C. All of CARME is >100 C except "Magnet Pipe" that is at 95 (probably OK given proximity to gate valve). 

13:40 - Decreased channel 5 to 160 C. Bottom back of CARME was going up, but rest of chamber was not following along. Thermal gradients were getting excessive.

15:10 - Checked Kapton wire specs from Allectra. Max temperature of 311-KAPM-060 is 260 C in a vacuum not 160 C as erroneously remembered. Increased Ch 1 2 3 to 160 and channel 5 to 170. Planning on leaving until weekend.


 

Progress can be followed on

http://web-docs.gsi.de/~lestinsk/baking.php

http://web-docs.gsi.de/~lestinsk/vacplot.php

Attachment 1: IMG_5600.JPG
IMG_5600.JPG
Attachment 2: IMG_5602.JPG
IMG_5602.JPG
Attachment 3: 20211117_122601.jpg
20211117_122601.jpg
Attachment 4: 20211117_151041.jpg
20211117_151041.jpg
Attachment 5: 20211117_151500.jpg
20211117_151500.jpg
Attachment 6: Kaptonwires-overview.pdf
Kaptonwires-overview.pdf Kaptonwires-overview.pdf Kaptonwires-overview.pdf Kaptonwires-overview.pdf
Attachment 7: CARMEtemps.png
CARMEtemps.png
  155   Tue Nov 16 12:05:13 2021 CB RSSCARME baking with tape continues

CARME baking with heating tape continues

Left overnight at 100C. Turned up this morning at 9:00 to 120C.

All good except detector strain relief seems to be going much higher than detector (no obvious reason) and then cool down. Proximity to flange / heating tape likely. One potential solution would be to push the motor in, putting some distance between detector / strain relief and outside flange.

 

13:30 - Increased channel 1 to 130 (long tape) and channels 5 and 6 to 140 (bottom flange heaters). Left rest as was.

15:20 - Increased channel 2 and 3 to 130. No obvious effects except Magnet pipe goes up

16:00 - Increased channel 1, 2, 3 to 140. Increased channels 5 and 6 to 150 C.

 

Progress can be followed on

http://web-docs.gsi.de/~lestinsk/baking.php

http://web-docs.gsi.de/~lestinsk/vacplot.php

Attachment 1: CARMEtemps.png
CARMEtemps.png
Attachment 2: 20211116_132048.jpg
20211116_132048.jpg
Attachment 3: 20211116_155138.jpg
20211116_155138.jpg
Attachment 4: 20211116_155857.jpg
20211116_155857.jpg
  154   Fri Nov 12 08:54:11 2021 RSS ML + CB (remotely) NK AK NPFirst CARME baking attempt

07:50 Small leaks (near the motors and upper valve) were detected in the tent which were covered with Al foil.

08:10 ML changed the baking temperature to 120 degrees.

09:40 Since the bottom of CARME is cold, due to uneven distribution of heat, baking is stopped.

13:00 Tent was slightly opened from the sides so that the heat can escape.

13:15 NK mounted three 200 W (see red labels in the picture below) and one 40 W (see the yellow label in the picture below) heating jacket on four CF100 flanges. Power cables and thermocouples were taken care of by AK.

14:00 Tent was closed.

14:00-17:00 ML started baking at temperature 80 degrees to check the heat distribution in the tent. Baking was then stopped and will start from Monday.

Attached - thermocouple temperatures (now correct)

Attachment 1: Capture.PNG
Capture.PNG
Attachment 2: 200W_heaters.svg
200W_heaters.svg
Attachment 3: 40W_heater.svg
40W_heater.svg
  153   Thu Nov 11 14:01:00 2021 AK, MLHeater channel layouts

Schematic layout of heaters and channel assignment attached. Table kindly provided by AK. I hope the conversion from excel to pdf didn't swallow any important information.

Attachment 1: 2020.11.05_Cryring-YR09-baking-layout.pdf
2020.11.05_Cryring-YR09-baking-layout.pdf
  151   Thu Nov 11 11:44:58 2021 RSS + ML OG + CB (remotely)CARME baking started

With the help of OG, the baking tent was built. MG started baking around 11:30 am with a rate of 0.2 K/minute.

Heating update can be found on:  https://web-docs.gsi.de/~lestinsk/baking.php (CARME temperature plot now visible as well - thanks Michael!)

where, device - 192.168.207.192 is for CARME.

Ch1: CARME chamber long tape (2 kW)

Ch2: CF100 flanges exit side

Ch3: CF100 bellows exit side

Ch5:  200 W (x3) CF 100 flanges

Ch6:  40 W CF 100 flange

Ch8: VAT 100 gate valve to ring

example: in 39:50, 39 is the temperature and 50 is the set temperature.

Attachment 1: IMG_5535.JPG
IMG_5535.JPG
Attachment 2: IMG_5537.JPG
IMG_5537.JPG
Attachment 3: IMG_5531.JPG
IMG_5531.JPG
Attachment 4: IMG_5534.JPG
IMG_5534.JPG
Attachment 5: IMG_5539.JPG
IMG_5539.JPG
  150   Wed Nov 10 18:38:16 2021 RSS TD + ML OG AK NKCARME wrapped in heating tape

Heating and electronic experts were in CRYRING and were unhappy with the damaged wire. Heaters were opened by OG, but no damage was found. It is suggested to replace the entire wire and then get it approved by the concerned authority.

To avoid further delays, ML and TD agreed on heating tapes/wires covering the chamber. OG helped in wrapping the heating wire around. Tent is to be built by tomorrow.

MG and RSS tried the magnetic power back off thing which eventually did not work.

AK and NK helped with the power and thermocouple of the heating wire. NK has also covered the pipe and bellows near the magnet with heating jackets.

 

Attachment 1: IMG_5514(1).JPG
IMG_5514(1).JPG
Attachment 2: IMG_5517(1).JPG
IMG_5517(1).JPG
Attachment 3: IMG_5518.JPG
IMG_5518.JPG
Attachment 4: IMG_5520.JPG
IMG_5520.JPG
Attachment 5: IMG_5521(1).JPG
IMG_5521(1).JPG
  149   Wed Nov 10 18:24:12 2021 RSSThermocouple labelling

1. Detector

2. Chamber back

3. Strain relief

4. Target

5. Front CARME

6. Top hanging

7. Middle middle CARME

8. Middle bottom

1. Middle top

2. End middle

3. Elbow

4. Valve

5. Magnet pipe

6. End back

7. Float

8. End bottom

Attachment 1: labeling.JPG.svg
labeling.JPG.svg
Attachment 2: Labelling.svg
Labelling.svg
Attachment 3: IMG_5512.JPG
IMG_5512.JPG
  147   Wed Nov 10 14:41:12 2021 TD RSS + CB (remotely)Wrapping CARME in heating tape
ML reports that baking of the upstream section (gas jet interaction chamber plus e-target) has commenced - initially to 50 deg C to check thermocouples. If OK baking 
temperature will be ramped slowly (which I think means c. 10 deg C per h) to limits established for specific components by the e-target.

Ragan and Oemer are currently threading a 10m heating tape helically around the CARME chamber.

16.20 ML reports baking of upstream section is proceeding OK - currently ramping to 100 deg C at 10 deg C / h.

16.25 RSS reports that he completed the install of the thermocouples this morning and will update the Elog with specific locations of each thermocuple and TC08 channel 
later today.

16:30 All CARME thermocouples <20 C. No temperature changes. Monitor.x left under screen. Please summon with screen -x and dismiss with CTRL+a d

17:40 One CARME thermocouple at 40C (CARME front)
  146   Tue Nov 9 15:03:29 2021 RSS + OGBurnt cables issues

Upper part of the frame was mounted. In order to do so, the valve was closed, turbo wire and hose were disconnected and then connected. Valve was opened again. See pictures 1-2.

Bottom part of the tent was constructed with 3 heaters (from one of the controllers) at the front, 4 in the middle, and 4 at the back (4+4 from the other controller). See pictures 3.

Thermocouples were tested and connected to TC08 modules.

Wire connecting to four heaters was damaged which caused a spark and sound when the upper tent was being constructed. This caused a power failure two times, turning off all three pumps. The upper valve was immediately closed. Power supplies of heaters were disconnected and pumps were turned on again. Valve was opened again as well.

Construction of tent is paused. Waiting for the wire to be repaired by tomorrow and start again.

Attachment 1: IMG_5493.JPG
IMG_5493.JPG
Attachment 2: IMG_5496.JPG
IMG_5496.JPG
Attachment 3: IMG_5497.JPG
IMG_5497.JPG
Attachment 4: IMG_5500.JPG
IMG_5500.JPG
Attachment 5: IMG_5502.JPG
IMG_5502.JPG
Attachment 6: IMG_5503.JPG
IMG_5503.JPG
Attachment 7: IMG_5507.JPG
IMG_5507.JPG
  145   Mon Nov 8 17:19:02 2021 RSS + OGNeg and ion cables connected

Neg and ion pump cables were labeled and connected.

Kapton tape wrapped around the neg cable connectors for the strain support.

Lower frame mounted.

Attachment 1: IMG_5480.JPG
IMG_5480.JPG
Attachment 2: IMG_5483.JPG
IMG_5483.JPG
Attachment 3: IMG_5488.JPG
IMG_5488.JPG
  144   Sat Nov 6 14:21:10 2021 CB RSSPreparation for bakeout tent mounting + to do for next week

Connected pneumatic hoses prepared by Gleb. Opened flow to pressure reducer set at 2 bar. Ok. Attempted to operate pneumatic motors. Somehow left / right pneumatic motors controls swapped around! Works fine otherwise. Most likely hardware issue, i.e. hoses are inverted inside motor box. Software fix possible in principle. Probably easiest to swap hoses going out from motor box (top to bottom and vice versa). Could not pull out hoses after installing. To do for next week.

 

Started connecting NEG and ion pumps. Realised wires are only labeled on one side. Had to separate all wires in one harness to make absolutely sure to know which wire is which. Second harness to be done next week. Please connect D2000 elements to bottom C4 power module, NEG getters wires 1-4 to middle power module and 5-6 to top power module. This means activation can proceed bottom to top of rack with minimal risk of mix ups. Please install top C4 module back in rack - missing appropriate rack shelf.

 

Installed Jan's Nim bin in rack. Jan thought that it was dead but seems to work just fine. Maybe test again?

 

 

To do next week 

Place thermocouples as per suggestion on file in baking Pi Baking directory.

Remove valves support. Install bottom of tent.

Install aluminium frame and heaters. This may require stopping the turbo pump and closing top gate valve.

Install rest of tent. Be careful not to pull or stretch the fabric. It's fragile to tension and fibreglass will pour out.

Install aluminised fibreglass and aluminium shell.

Start baking. Target temperature 120C.

  143   Fri Nov 5 12:27:11 2021 CB RSS + DR NP ML USFirst pump down of YR09

Tightened all flanges.

Added Pirani gauge and third IE514 to interaction chamber.

Stated pump down. Around 100 mbar on Interaction chamber Pirani gauge very loud metallic noise is heard. Found bellows between CARME and interaction chamber completely pressed flat. CARME and interaction chamber now much closer. Vented immediately. No pictures. Bellows return to normal state.

 

Repeated slowly and carefully using theodolite laser on reference point on CARME. Found CARME does not move appreciably forward, thus motion must come from interaction chamber frame.

 

Mounted two feet between CARME and interaction chamber. Pumped down again. Interaction chamber now moves by few mm. No issues with bellows. Left pumping overnight (Thu 4 November)

 

Found at 1E-6 mbar this morning. Leak tested. Found four minor leaks. Fixed by tightening. 

 

US + Kosmos installed Be window. Pumping down extremely slowly. 1 mbar per 2 second. Started turbo at 1 mbar. Pumping down overnight. No damage to window.

 

See attached new plastic free struts and motion of interaction chamber before after pumping. See skull nose.

Attachment 1: 20211105_131748.jpg
20211105_131748.jpg
Attachment 2: 20211105_135918.jpg
20211105_135918.jpg
Attachment 3: 20211105_135929.jpg
20211105_135929.jpg
Attachment 4: 20211105_140451.jpg
20211105_140451.jpg
Attachment 5: 20211105_150616.jpg
20211105_150616.jpg
  142   Wed Nov 3 16:39:01 2021 CB RSS + NP DR NK MLInteraction chamber lifted in place

Prepared to move interaction chamber in place. Ok CARME side, flange has M10 bolts. On interaction chamber side i.e. bellows it has M8 bolts.

DS designed this to have loose fit under assumption of  M10 side actually being 3/8 holes i.e. 9.9 mm. However holes are 10.2 mm. Furthermore Grossmontage suggest loose fit will be damaged in baking in the long term.

 

Dismounted bellow ring from spare bellows. Re-drilled to M10 and cleaned. Remounted. Old bellow outer ring moved with spare bellows.

 

Lifted interaction chamber in place. Alignment started. Will finish tomorrow morning, followed by pump down and leak test. Last step before beginning final pump down should be mounting Be window on Friday. Need US for that.

Attachment 1: 20211103_161553.jpg
20211103_161553.jpg
Attachment 2: 20211103_161813.jpg
20211103_161813.jpg
Attachment 3: 20211103_175457.jpg
20211103_175457.jpg
  141   Wed Nov 3 11:18:11 2021 CB RSS + DR NK MLLeak closed

Found CARME at 1.0E-6 mbar. Beginning leak test.

 

Found leaks in the same location as before. Roughly 1~3E-8 mbar l second vs. base leak of 1E-12

 

Tightened right hand half of bolts to 120 Nm. No changes but leak goes up to 1E-10 mbar l second when tightening bolts closer to leak location, without He being sprayed.

Tightened bolts closer to leak to 140. Leak goes down to 1E-9 mbar l second

Tightened four bolts closest to leak to 170 Nm. Leak disappeared (<1E-11 mbar). Tested repeatedly all around CARME including left hand side of 600 COF flange. No leaks.

 

Vented. Installing jet target in the afternoon.

 

 

Attachment 1: 20211103_110244.jpg
20211103_110244.jpg
Attachment 2: 20211103_110414.jpg
20211103_110414.jpg
  140   Tue Nov 2 12:45:36 2021 CB + DR NKIssues with CARME feet

Attempted to close CARME as carefully as possible.

Swapped out one foot for another one and temporarily removed a second to allow CARME to rotate slightly and get a better alignment of the two surfaces of the COF600 flange.

 

Closed CARME but realised carriages with struts are such that carriages touch before flanges do. Worse on side with leak than side without. See attached. As a result we would not be able to close properly.

 

DR to modify the holes in the feet closest to flange on CARME front to allow surfaces to touch before carriages come in contact. Hopefully close later today.

 

Issue was not encountered earlier because CARME was fully mounted before being dropped on feet. Struts were forcefully brought close to allow CARME to descend. Unclear how this was done in DL but may be baking had altered some sizes enough for the issue to appear. Spare feet do not help significantly.

 

Closed CARME . 1. Bolts finger tight. 2. Bolts tightened with tools stopping as soon as resistance is felt. Tightening done from the top descending towards the right (clockwise) and from the bottom counterclockwise. 3. 20 Nm applied in the same way 3. 40 4. 60 5. 80 6 80 again but single operator going round 6. 100 single operator

 

Pumping down overnight

 

 

 

 

 

Attachment 1: 20211102_134446.jpg
20211102_134446.jpg
  139   Mon Nov 1 14:48:16 2021 CB + ML DRCARME opened

Confirmed leak as per previous Elog.

Tightened to 100 Nm - no effects.

Tightened to 130 Nm - bottom leak disappeared. Top leak still at 1.3E-8 mbar l second.

 

Decided to open. Lifted front part of CARME down. 

No obvious damage to wire seal.

Inspected wire seat and seal part. Part corresponding to leak feels rough to the touch. Suggested damage when SS parts come together? 

DR sanded down gently the valve seal. Cleaned with compressed air and ethanol. 

Gently sanded down seat as well, barely touched it.

 

Grossmontage experts will be back tomorrow am for final cleaning and re-assembly.

  138   Sun Oct 31 13:55:13 2021 CB OHLeak tests

Found CARME at 2E-7.

Closed valve on top of roughing pump. Clamp was not tight. Pressure still good.

 

Connected to leak cart instead. Brought leak cart to 1E-2. Did not open valve yet. Found CARME pressure at 7.5 mbar. Leak in the backing line? Unclear what happened but moving backing line flexible hose increases pressure. Tightened all clamps. Moving backing flexible hose no longer seems to have effect.

 

Pumping down again for leak tests. Turbo looks and sounds fine.

 

Recovered E-6 in few minutes. Reconnected leak cart. Pressure does not increase. Still not sure what happened earlier. Beginning leak tests.

 

Leaks found in the wire seal around location of picture attached. Would suggest tightening wire seal tomorrow and check again. Leak must be small given base pressure achieved.

 

 

Attachment 1: 20211031_155820.jpg
20211031_155820.jpg
  137   Sun Oct 31 08:05:15 2021 CB OH + DR MLCARME closed again + feedthroughs labeled

(Fri 29 October)

Cleaned wire seal recess and labeled thermocouples and flanges to identify p-n vs. n-n side and ports with bias pin. See attached in case baking damages the label.

Grossmontage insisted to clean wire seal once more to remove black stains noticed when opening. Kapton vapour condensation could be possible cause. They used special cleaning product from XHV (not sure what - acetone mix with something?)

Lifted CARME back. Inserted wire seal. Closed chamber. Step is noted between central CARME section and other two - it's normal. Noted some damage to SS vessel on the outside. Unclear what could have caused it but it's only cosmetic.

Wire seal tightened with 20 - 40 - 60 - 80 - 80 Nm turns starting from top and proceeding in a circle. Had to temporarily dismount on foot to make space for torque wrench. Also had to saw off front part of ratchet attachmnent to slide wrench between GSI ion pump and flange. All other flanges closed by Grossmontage. Pump down started.

Attachment 1: 20211029_110228.jpg
20211029_110228.jpg
Attachment 2: 20211029_110617.jpg
20211029_110617.jpg
Attachment 3: 20211029_115220.jpg
20211029_115220.jpg
Attachment 4: 20211029_130422.jpg
20211029_130422.jpg
Attachment 5: LowerFeedthroughs.png
LowerFeedthroughs.png
Attachment 6: PXL_20211029_074516697.jpg
PXL_20211029_074516697.jpg
Attachment 7: PXL_20211029_080123618.jpg
PXL_20211029_080123618.jpg
Attachment 8: PXL_20211029_080130072.jpg
PXL_20211029_080130072.jpg
Attachment 9: PXL_20211029_080130885.jpg
PXL_20211029_080130885.jpg
Attachment 10: PXL_20211029_080136793.jpg
PXL_20211029_080136793.jpg
Attachment 11: PXL_20211029_080139929.jpg
PXL_20211029_080139929.jpg
  136   Sun Oct 31 07:55:39 2021 CB JM + OHStrain relief mounted

(Thu 28 October)

All strain relief extension plates mounted behind the wires, using SS countersunk bolts (not silver plated).

Inserting bottom part of MACOR strain relief causes wires to bend too much with strain placed on DSSD. No reasonable way to proceed like that.

Opted to move hard kapton coated copper wire to line up with holes in top MACOR strain relief, and clamp the wrapped wire bundles between top MACOR strain relief and extension plate on bottom.

Results good. No significant movement when wires are moved or motors actuated. See videos - password abcd1234


https://datasync.ed.ac.uk/index.php/s/5waHTSY9husEgy9

https://datasync.ed.ac.uk/index.php/s/34293z7kQs2Mppd

https://datasync.ed.ac.uk/index.php/s/xmHFVqqIchtixGT
 

 

Mounted plate with MACOR shield on other rod. Connected thermocouples. Re-mounted 2x UHV1400 getters.

 

 

With help from OH attempted V-I test of detectors. SHV to 2x lemo broken (faulty lemo connector). Had to borrow AID@DESPEC cable.

Diode function OK. VI tests OK but inconclusive. Both DSSDs drain current when voltage is applied, but current not quite stable. Light leak very likely. Will attempt again after closing.

Attachment 1: 20211028_190026.jpg
20211028_190026.jpg
Attachment 2: 20211028_192251.jpg
20211028_192251.jpg
Attachment 3: 20211028_192210.jpg
20211028_192210.jpg
ELOG V3.1.4-unknown